A Young’s double slit apparatus has slits separated by 0.28 mm and a screen 48 cm away from the slits. The whole apparatus is immersed in water and the slits are illuminated by the red light (λ = 700 nm in vacuum). Find the fringe-width of the pattern formed on the screen.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. Given that, d = 0.28 mm = 0.28 * 10^–3 m, D = 48 cm = 0.48 m, λ base a = 700 nm in vacuum Let, λ base w = wavelength of red light in water Since, the fringe width of the pattern is given by, Β = λ base wD/d = 525 * 10^-9 * 0.48/0.28 * 10^-3 = 9 * 10^-4 m = 0.90 mm.